OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 100. Find y when x = 6

Parth (parthkohli):

So basically: \(\Rightarrow y = kx^2\)

Parth (parthkohli):

You can find k by substitution. \(\Rightarrow 100 = k \times 25 \) \(\Rightarrow k = 4\)

OpenStudy (anonymous):

I believe Y would = 120..?

Parth (parthkohli):

Now apply the same formula \(y = kx^2\) when x = 6 and k = 4 \(\Rightarrow y = 4 \times 6^2 \) \(\Rightarrow y = 4 \times 36\) \(\Rightarrow y = 144\)

Parth (parthkohli):

No....

OpenStudy (anonymous):

Okay, I understand it.

Parth (parthkohli):

Excellent!

OpenStudy (anonymous):

i believe the answer is y=144 XD..